Eight guests have to be seated 4 on each sides of a long rectangular table 2 particular guests desire to sit on one side of the table and 3 on the other . The number of ways in which the sitting arrangements can be made is ?
Answer by addingup(3677)

The table has two sides, select one: = 2
Arrange 2 on one side: 2! = 2 (1*2)
Arrange 3 on the other side: 3! = 6 (1*2*3)
We now have 3 guests that still need to be seated.
Select 2 out of 3: 3 choose 2 = 3
Place them on the side with 2: 1
Place the remaining guest on the side with 3: 1
--------------------
Total possible sitting arrangements: 2*2*6*3*1*1 = 72